OpenStudy (luvhawks711):

would these make a right triangle? explain your answer. a. 15 feet, 36 feet, and 39 feet b. 20 inches, 21 inches, and 29 inches c. 8 yards, 9 yards, and 12 yards d. 4 meters, 7 meters, and 8 meters

OpenStudy (luvhawks711):

@Abhisar

OpenStudy (luvhawks711):

@charliemath

OpenStudy (welshfella):

right angled triangles obey the pythagoras theorem a^2 = b^2 + c^2 where a is the hypotenuse ( the longest side)

OpenStudy (welshfella):

so it 39^2 = 36^2 + 15^2 then a is right angled

OpenStudy (welshfella):

* so if i meant to write

OpenStudy (luvhawks711):

what about (b), (c), (d) @welshfella

OpenStudy (welshfella):

the same thing applies to them

OpenStudy (welshfella):

if they dont fit the formula then they are not right-angled

OpenStudy (luvhawks711):

ok

OpenStudy (welshfella):

39^2 = 1521 36^2 = 1296 15^2 = 225 so what do you think about (a) ?

OpenStudy (welshfella):

add the last 2 up

OpenStudy (welshfella):

1296+225 = ?

OpenStudy (welshfella):

grab your calculator and chug away...